OpenStudy (anonymous):

Debbie begins working on a test at 12:04 and finished the test a few minutes later at 12:26. By how many degrees has the distance between the minute hand and the hour hand increased during this time interval? Assume the smallest angle formed in your calculations.please show your work

OpenStudy (anonymous):

The minute hand passes through \(\frac{1}{60}\) of a turn every minute. The hour hand passes through \(\frac{1}{12}\times\frac{1}{60}\) of a turn every minute.